what is the lowest common mutltiple of 8,28,48
Answer by orca(409)

STEP 1
First factor each integer into primes:
8 = 2*2*2
28 = 2*2*7
48 = 2*2*2*2*3
STEP 2
Next find the maximum number of time each prime appears in each factorization.
The maximum number of time the prime "2" appears is 4 in the factorization of 48.
The maximum number of time the prime "3" appears is 1 in the factorization of 48.
The maximum number of time the prime "7" appears is 1 in the factorization of 28.
STEP 3
Finally, the lowest common multiple of 8,28,48 can be formed by writing each prime the maximum number of times
So write 2 four times, 3 one time and 7 one time.
Thus the lowest common multiple of 8,28,48 is 2*2*2*2*3*7 = 336.